Perform the indicated operation. Write the result in simplest form: 2 9/10 + 5 2/3 - 1 1/6.
I am unsure of the result of this problem. Please explain. Thanks.

Answer by checkley79(3341)

2 9/10 + 5 2/3 - 1 1/6.
(2*10+9)/10+(5*3+2)/3-(1*6+1)/6 Now find a common denominator.
29/10+17/3-7/6
10=1*2*5
3=1*3
6=1*2*3
Elements are: 1*2*3*5=30.
(3*29+10*17-7*5)/30
(87+170-35)/30
222/30=37/5=7 2/5 ans.